OpenStudy (cenaida):

What is the common ratio of this geometric sequence? 1 80 2 40 3 20 4 10 5 5 A. -40 B. 0.5 C. 2 D. 40

OpenStudy (***[isuru]***):

hi,\[common \ ratio (r) = \frac{ second \ term }{ first \ term }\] wt will be ur answer then ?

OpenStudy (cenaida):

D?

OpenStudy (***[isuru]***):

\[\frac{ second \ term }{ first \ term } = \frac{ 40 }{80 }\] Did u get it now?

OpenStudy (***[isuru]***):

now.... can u go for the answer ?

OpenStudy (cenaida):

oooh so its C ?

OpenStudy (***[isuru]***):

nope...dr \[\frac{ 40 }{ 80 } =\frac{ 4 }{8 } = \frac{ 1 }{ 2}\] It cant be C

OpenStudy (anonymous):

its 0.5

OpenStudy (cenaida):

Thanks!

OpenStudy (***[isuru]***):

u r welcome!!